OpenStudy (hhlppleasevv):

PLEASE HELP An initial population of 745 quail increases at an annual rate of 16%. Write an exponential function to model the quail population. What will the approximate population be after 4 years?

OpenStudy (wolf1728):

y(t) = a × (e)^ k*t y(t) = final amount a = beginning amount k = rate t=time e=2.718281828 ending amount = 745 * 2.718281828^.16*4 ending amount = 745 * 2.718281828^.64